# Sampling from the Mean-Field Stationary Distribution

### v2

@@ B.1 Convexity and Smoothness
$\qquad+\frac{1}{N-1}\,\underbrace{\begin{bmatrix}\sum_{j=2}^{N}\nabla^{2}W(x^{1}-x^{j})&-\nabla^{2}W(x^{1}-x^{2})&\cdots&-\nabla^{2}W(x^{1}-x^{N})\\ -\nabla^{2}W(x^{2}-x^{1})&\sum_{\begin{subarray}{c}j=1\\ j\neq 2\end{subarray}}^{N}\nabla^{2}W(x^{2}-x^{j})&\cdots&-\nabla^{2}W(x^{2}-x^{N})\\ \vdots&\vdots&\ddots&\vdots\\ -\nabla^{2}W(x^{N}-x^{1})&-\nabla^{2}W(x^{N}-x^{2})&\cdots&\sum_{j=1}^{N-1}\nabla^{2}W(x^{N}-x^{j})\end{bmatrix}}_{=\mathsf{B}}\,.$
@@ B.2 Bounded Perturbations
We now prove the statement about $\mu^{1:N}$. By the tensorization argument in [BGL14, Proposition 5.2.7] or [BLM13, Theorem 4.10], we obtain the inequality that for any smooth test function $g^{2}:\mathbb{R}^{d\times N}\to\mathbb{R}_{+}$ and $X^{-i}\sim\mu^{-i}$,
$\operatorname{\mathsf{ent}}_{\mu^{1:N}}(g^{2})\leq\sum_{i=1}^{N}\operatorname{\mathbb{E}}\operatorname{\mathsf{ent}}_{\mu^{i|-i}(\cdot|X^{-i})}\bigl{(}g^{2}(X^{1:i-1},\cdot,X^{i+1:N})\bigr{)}\,,$ (B.1)
where each entropy term in RHS involves the distribution of the $i$-th particle, conditioned on the remaining particles. This conditional measure has a density of the form
@@
$C_{\mathsf{LSI}}\leq\frac{\sigma^{2}}{2}\,\bigl{(}\alpha_{V_{0}}+\frac{N}{N-1}\,{\alpha_{W_{0}}^{-}}\bigr{)}^{-1}\exp\Bigl{(}\frac{2}{\sigma^{2}}\,\bigl{(}\operatorname{\mathsf{osc}}(V_{1})+\operatorname{\mathsf{osc}}(V_{1})\bigr{)}\Bigr{)}\,.$
Therefore, we can further bound each entropy term in (B.1) by
$\operatorname{\mathsf{ent}}_{\mu^{1:N}}(g^{2})\leq\sum_{i=1}^{N}2C_{\mathsf{LSI}}\operatorname{\mathbb{E}}\operatorname{\mathbb{E}}_{Z\sim\mu^{i|-i}(\cdot\mid X^{-i})}[\lvert\partial_{i}g(X^{1:i-1},Z,X^{i+1:N})\rvert^{2}]=2C_{\mathsf{LSI}}\operatorname{\mathbb{E}}_{\mu^{1:N}}[\lVert\nabla g\rVert^{2}]\,.$
Therefore, $\mu^{1:N}$ satisfies (LSI) with parameter $C_{\mathsf{LSI}}$. ∎
@@ B.3 Logarithmic Sobolev Inequalities via Perturbations
Lemma 20 (LSI under Lipschitz Perturbations [CG22, Theorem 2.7]).
Let $\nu\propto\exp(-V)$ for an $\alpha_{V}$-strongly convex and $\beta_{V}$-smooth function $V:\mathbb{R}^{d}\to\mathbb{R}$. Let $H:\mathbb{R}^{d}\to\mathbb{R}$ be a $L$-Lipschitz continuous function. Then, $\mu\propto\exp(-H)\,\nu$ satisfies (LSI) with constant $C_{\mathsf{LSI}}(\mu)$ given by
$C_{\mathsf{LSI}}(\mu)\leq\frac{4}{\alpha_{V}}+\Bigl{(}\frac{L}{\alpha_{V}}+\sqrt{\frac{2}{\alpha_{V}}}\Bigr{)}^{2}\,\Bigl{(}2+d+\frac{d}{2}\log\frac{\beta_{V}}{\alpha_{V}}+\frac{4L^{2}}{\alpha_{V}}\Bigr{)}\exp\Bigl{(}\frac{L^{2}}{2\alpha_{V}}\Bigr{)}\,.$
@@
Lemma 21 (Uniform LSI for the Proximal Gibbs Measure [SWN23, Theorem 1]).
For the proximal Gibbs measure (2.3) in the setting of §2.1.2, under Assumptions 6, and 8, we have that
$\sup_{\mu\in\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})}C_{\mathsf{LSI}}(\pi_{\mu})\leq\overline{C}_{\mathsf{LSI}}\,,$
@@
$\overline{C}_{\mathsf{LSI}}\leq\frac{\sigma^{2}}{\beta}\exp\Bigl{(}\frac{8B^{2}}{\lambda\sigma^{2}}\sqrt{2d/\pi}\Bigr{)}\wedge\Bigl{\{}\frac{2\sigma^{2}}{\beta}+\bigl{(}\frac{B}{\beta}+\frac{\sigma}{\sqrt{\beta}}\bigr{)}^{2}\,\bigl{(}2+d+\frac{8B^{2}}{\beta\sigma^{2}}\bigr{)}\exp\bigl{(}\frac{B^{2}}{\lambda\sigma^{2}}\bigr{)}\Bigr{\}}\,.$
Finally, we provide an LSI for the finite-particle stationary distribution $\mu^{1:N}$ with a constant independent of $N$, using an approximate tensorization argument.
Lemma 22.
Consider the general McKean–Vlasov setting with $\mathcal{F}(\mu)=\mathcal{F}_{0}(\mu)+\frac{\lambda}{2}\int\lVert\cdot\rVert^{2}\,\mathrm{d}\mu$ and suppose that Assumption 8 holds. Then, $\mu^{1:N}$ satisfies (LSI) with parameter independent of $N$ given by
@@
$C_{\mathsf{LSI}}(\mu^{1:N})\leq\frac{4\sigma^{2}}{\lambda}+\Bigl{(}\frac{2B}{\lambda}+\frac{\sigma}{\sqrt{\lambda}}\Bigr{)}^{2}\,\Bigl{(}2+d+\frac{16B^{2}}{\lambda\sigma^{2}}\Bigr{)}\exp\Bigl{(}\frac{2B^{2}}{\lambda\sigma^{2}}\Bigr{)}\,.$
Proof.  Since $\mu^{1:N}(x^{1:N})\propto\exp(-\frac{2N}{\sigma^{2}}\,\mathcal{F}(\rho_{x^{1:N}}))$, where $\rho_{x^{1:N}}=\frac{1}{N}\sum_{i=1}^{N}\delta_{x^{i}}$ is the empirical measure, we can check that
@@
$\mu^{i\mid-i}(x^{i}\mid x^{-i})\propto\exp\Bigl{(}-\frac{2N}{\sigma^{2}}\,\mathcal{F}_{0}(\rho_{x^{1:N}})-\frac{\lambda}{\sigma^{2}}\,\lVert x^{i}\rVert^{2}\Bigr{)}\,.$
Also, note that
$\nabla_{x^{i}}\mathcal{F}_{0}(\rho_{x^{1:N}})$ $=\frac{1}{N}\,\nabla_{\mathcal{W}_{2}}\mathcal{F}_{0}(\rho_{x^{1:N}},x^{i})\,.$
Using Assumption 8, $x^{i}\mapsto\frac{2N}{\sigma^{2}}\,\mathcal{F}_{0}(\rho_{x^{1:N}})$ is $\nicefrac{{2B}}{{\sigma^{2}}}$-Lipschitz.
Then, apply Lemma 20 with $\alpha_{V}=\beta_{V}=\nicefrac{{\lambda}}{{\sigma^{2}}}$, to obtain

### v4

@@ B.1 Convexity and Smoothness
$\quad+\frac{1}{N-1}\,\underbrace{\begin{bmatrix}\sum_{j=2}^{N}\nabla^{2}W(x^{1}-x^{j})&-\nabla^{2}W(x^{1}-x^{2})&\cdots&-\nabla^{2}W(x^{1}-x^{N})\\ -\nabla^{2}W(x^{2}-x^{1})&\sum_{\begin{subarray}{c}j=1\\ j\neq 2\end{subarray}}^{N}\nabla^{2}W(x^{2}-x^{j})&\cdots&-\nabla^{2}W(x^{2}-x^{N})\\ \vdots&\vdots&\ddots&\vdots\\ -\nabla^{2}W(x^{N}-x^{1})&-\nabla^{2}W(x^{N}-x^{2})&\cdots&\sum_{j=1}^{N-1}\nabla^{2}W(x^{N}-x^{j})\end{bmatrix}}_{=\mathsf{B}}\,.$
@@ B.2 Bounded Perturbations
Before proceeding to the proof of Lemma 8, we first state a result on log-Sobolev inequalities under weak interactions due to [otto2007new].
Lemma 20 ([otto2007new, Theorem 1]).
Consider a measure $\mu^{1:N}$ on $\mathbb{R}^{d\times N}$, with conditional measures $\mu^{i|-i}$. Assume that
$C_{\mathsf{LSI}}(\mu^{i|-i}(\cdot\mid x^{-i}))$ $\leq\frac{1}{\tau_{i}}\,,\qquad\text{for all $i\in[N]$, $x^{-i}\in\mathbb{R}^{d\times(N-1)}$}\,,$
$\lVert\nabla_{x^{i}}\nabla_{x^{j}}\log\mu^{1:N}(x^{1:N})\rVert$ $\leq\beta_{i,j}\,,\qquad\text{for all $x^{1:N}\in\mathbb{R}^{d\times N}$, $i,j\in[N]$, $i\neq j$}\,.$
Then, consider the matrix $A\in\mathbb{R}^{N\times N}$ with entries $A_{i,i}=\tau_{i}$, $A_{i,j}=-\beta_{i,j}$ for $i\neq j$. If $A\succeq\rho I_{N}$, then $\mu^{1:N}$ satisfies (LSI) with constant $1/\rho$.
@@
We now prove the statement about $\mu^{1:N}$. Each conditional measure has a density of the form
@@
$\tau_{i}^{-1}\leq\tau^{-1}\coloneqq\frac{\sigma^{2}}{2}\,\bigl{(}\alpha_{V_{0}}+\frac{N}{N-1}\,{\alpha_{W_{0}}^{-}}\bigr{)}^{-1}\exp\Bigl{(}\frac{2}{\sigma^{2}}\,\bigl{(}\operatorname{\mathsf{osc}}(V_{1})+\operatorname{\mathsf{osc}}(V_{1})\bigr{)}\Bigr{)}\,.$
Secondly, we note that from (2.1), we have
$\frac{2}{\sigma^{2}\,(N-1)}\sup_{z\in\mathbb{R}^{d}}\lVert\nabla^{2}W(z)\rVert_{\mathsf{op}}\leq\frac{2\beta_{W}}{\sigma^{2}\,(N-1)}\eqqcolon\beta_{i,j}\,.$
Thus, we have
$A=\begin{bmatrix}\tau&-\frac{2\beta_{W}}{\sigma^{2}\,(N-1)}&\cdots&-\frac{2\beta_{W}}{\sigma^{2}\,(N-1)}\\[5.0pt] -\frac{2\beta_{W}}{\sigma^{2}\,(N-1)}&\tau&\cdots&-\frac{2\beta_{W}}{\sigma^{2}\,(N-1)}\\[5.0pt] \vdots&\vdots&\ddots&\vdots\\[5.0pt] -\frac{2\beta_{W}}{\sigma^{2}\,(N-1)}&-\frac{2\beta_{W}}{\sigma^{2}\,(N-1)}&\cdots&\tau\end{bmatrix}\,.$
Under Assumption 10, this matrix is strictly diagonally dominant and has a minimum eigenvalue of at least $\tau/2$. We can now apply Lemma 20 to complete the proof. ∎
@@ B.3 Logarithmic Sobolev Inequalities via Perturbations
Lemma 21 (LSI under Lipschitz Perturbations [BriPed24HeatFlow, Theorem 1.4]).
Let $\mu\propto\exp(-H-V)$ for an $\alpha_{V}$-strongly convex function $V:\mathbb{R}^{d}\to\mathbb{R}$ and an $L$-Lipschitz function $H:\mathbb{R}^{d}\to\mathbb{R}$. Then, $\mu$ satisfies (LSI) with constant $C_{\mathsf{LSI}}(\mu)$ given by
$C_{\mathsf{LSI}}(\mu)\leq\frac{1}{\alpha}\exp\Bigl{(}\frac{L^{2}}{\alpha}+\frac{4L}{\sqrt{\alpha}}\Bigr{)}\,.$
@@
Lemma 22 (Uniform LSI for the Proximal Gibbs Measure [suzuki2023convergence, Theorem 1]).
For the proximal Gibbs measure (2.3) in the setting of §2.1.2, under Assumption 8, we have that
$\sup_{\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})}C_{\mathsf{LSI}}(\pi_{\mu})\leq\overline{C}_{\mathsf{LSI}}\,,$
@@
$\overline{C}_{\mathsf{LSI}}$ $\leq\frac{\sigma^{2}}{2\lambda}\exp\Bigl{(}\frac{2B^{2}}{\lambda\sigma^{2}}+\frac{8B}{\sqrt{2\lambda}\,\sigma}\Bigr{)}\,.$
Obtaining a uniform-in-$N$ LSI for $\mu^{1:N}$ under Assumption 8 is more difficult, and we rely on the recent heat flow estimates of [BriPed24HeatFlow]. In their work, the authors showed the existence of an $L$-Lipschitz transport map—the Kim–Milman map [KimMil12ReverseHeat]—from the standard Gaussian measure $\gamma$ to a measure $\mu$, under suitable assumptions on $\mu$. By [bakry2014analysis, Proposition 5.4.3], this immediately implies that $\mu$ satisfies $C_{\mathsf{LSI}}(\mu)\leq L^{2}$. The existence of the Lipschitz transport map is based on estimates along the heat flow, and we summarize the computation in a convenient form based on bounding the operator norm of the covariance matrix of Gaussian tilts of the measure. The latter property is sometimes called tilt stability in the literature. Note that we do not attempt to optimize constants here.
Lemma 23 (Lipschitz Transport Maps via Reverse OU).
Let $\mu$ be a probability measure over $\mathbb{R}^{d}$ and for any $t>0$, $y\in\mathbb{R}^{d}$, let $\mu_{t,y}$ denote the Gaussian tilt,
$\mu_{t,y}(\mathrm{d}x)\propto\exp\Bigl{(}-\frac{\lVert y-x\rVert^{2}}{2t}+\frac{\lVert x\rVert^{2}}{2}\Bigr{)}\,\mu(\mathrm{d}x)\,,$ (B.1)
where we assume that this defines a valid probability measure for all $t>0$ and $y\in\mathbb{R}^{d}$. Suppose there exist $a,C>0$ such that the following “tilt stability” property holds:
$\lVert\operatorname{cov}_{\mu_{t,y}}\rVert_{\mathsf{op}}$ $\leq\Bigl{(}\frac{1}{\sqrt{a+1/t}}+\frac{C}{a+1/t}\Bigr{)}^{2}\,,\qquad\text{for all}~{}t>0\,,\;y\in\mathbb{R}^{d}\,.$
Then, there exists an $L$-Lipschitz transport map $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $T_{\#}\gamma=\mu$, where $\gamma$ is the standard Gaussian measure and $L$ can be estimated by
$L$ $\leq\frac{1}{\sqrt{1+a}}\exp\Bigl{(}\frac{C^{2}}{2\,(1+a)}+\frac{2C}{\sqrt{1+a}}\Bigr{)}\,.$
Proof.  We follow the calculations of [BriPed24HeatFlow]. Let ${(P_{t})}_{t\geq 0}$ denote the heat semigroup, and ${(Q_{t})}_{t\geq 0}$ the Ornstein–Uhlenbeck semigroup. Then, if $\gamma$ denotes the standard Gaussian measure, the identity $Q_{t}f=P_{1-\exp(-2t)}f(\exp(-t)\,\cdot)$ and
$-\frac{I}{\exp(2t)-1}$ $\preceq\nabla^{2}\log Q_{t}\bigl{(}\frac{\mu}{\gamma}\bigr{)}=\exp(-2t)\,\Bigl{[}\nabla^{2}\log P_{1-\exp(-2t)}\bigl{(}\frac{\mu}{\gamma}\bigr{)}\Bigr{]}(\exp(-t)\,\cdot)$
$=\frac{1}{\exp(2t)-1}\,\Bigl{(}\frac{\operatorname{cov}_{\mu_{1-\exp(-2t),\,\exp(-t)\,\cdot}}}{1-\exp(-2t)}-I\Bigr{)}$
$\preceq\frac{I}{\exp(2t)-1}\,\Bigl{[}\Bigl{(}\frac{1/\sqrt{1-\exp(-2t)}}{\sqrt{a+1/(1-\exp(-2t))}}+\frac{C/\sqrt{1-\exp(-2t)}}{a+1/(1-\exp(-2t))}\Bigr{)}^{2}-1\Bigr{]}$
$\preceq\Bigl{[}\frac{1-\alpha}{\alpha\,(\exp(2t)-1)+1}+\frac{C^{2}\exp(2t)}{(\alpha\,(\exp(2t)-1)+1)^{2}}$
$\qquad\qquad\qquad\qquad\qquad{}+\frac{2C\exp(2t)}{(\exp(2t)-1)^{1/2}\,(\alpha\,(\exp(2t)-1)+1)^{3/2}}\Bigr{]}\,I$
where $a=\alpha-1$. Note that we can identify this with the bound of [BriPed24HeatFlow, Corollary 3.2] with $C=L$. Thus, by following the calculations in the proof of [BriPed24HeatFlow, Theorem 1.4], we obtain the desired result. ∎
We next verify the tilt stability condition, leveraging the propagation of chaos result in Theorem 5.
Lemma 24 (Tilt Stability).
For any $t>0$ and $y^{1:N}\in\mathbb{R}^{d\times N}$, let $\mu_{t,y^{1:N}}$ be the Gaussian tilt of $\mu^{1:N}$ as defined in (B.1) with $\mu^{1:N}$ in place of $\mu$. Here, $\mu^{1:N}$ is the stationary distribution (1.1) of the mean-field Langevin diffusion. Then, under Assumptions 56, and 8, and assuming that $2\lambda>\sigma^{2}$ and
$N\gtrsim\frac{\beta d}{2\lambda-\sigma^{2}}\exp\Bigl{(}\frac{8B^{2}/\sigma^{4}}{2\lambda/\sigma^{2}-1}\Bigr{)}$ (B.2)
for a sufficiently large implied (but universal) constant, we have
@@
$\lVert\operatorname{cov}_{\mu_{t,y^{1:N}}}\rVert_{\mathsf{op}}$ $\leq\Bigl{[}\frac{1}{\sqrt{2\lambda/\sigma^{2}-1+1/t}}+\frac{1}{2\lambda/\sigma^{2}-1+1/t}\,\mathcal{O}\Bigl{(}\frac{B}{\sigma^{2}}+\frac{\sqrt{\beta d}}{\sigma}\exp\frac{8B^{2}/\sigma^{4}}{2\lambda/\sigma^{2}-1}\Bigr{)}\Bigr{]}^{2}\,.$
Proof.  We introduce the auxiliary measures
$\pi_{t,y^{i}}(x^{i})\propto\exp\Bigl{(}-\frac{\lVert y^{i}-x^{i}\rVert^{2}}{2t}-\bigl{(}\frac{\lambda}{\sigma^{2}}-\frac{1}{2}\bigr{)}\,\lVert x^{i}\rVert^{2}-\frac{2}{\sigma^{2}}\,\delta\mathcal{F}_{0}(\bar{\pi}_{t,y^{1:N}},x^{i})\Bigr{)}\,,\quad i\in[N]\,,$ (B.3)
where $\bar{\pi}_{t,y^{1:N}}\coloneqq\frac{1}{N}\sum_{i=1}^{N}\pi_{t,y^{i}}$. To see that these auxiliary measures are well-defined, note that any minimizer of the functional
$(\pi^{1},\dotsc,\pi^{N})$